# When Sharpening Becomes Collapse: Sampling Bias and Semantic Coupling in RL with Verifiable Rewards

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) is a central paradigm for turning large language models (LLMs) into reliable problem solvers, especially in logic‑heavy domains. Despite its empirical success, it remains unclear whether RLVR elicits novel capabilities or merely sharpens the distribution over existing knowledge. We study this by formalizing over-sharpening, a phenomenon where the policy collapses onto limited modes, suppressing valid alternatives. At a high level, we discover finite-batch updates intrinsically bias learning toward sampled modes, triggering a collapse that propagates globally via semantic coupling. To mitigate this, we propose inverse-success advantage calibration to prioritize difficult queries and distribution-level calibration to diversify sampling via a memory network. Empirical evaluations validate that our strategies can effectively improve generalization.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a cornerstone approach for specializing large language models (LLMs) to domains where correctness is checkable (DeepSeek-AI, 2025; Yang et al., 2025). However, there is still no clear consensus on what RLVR fundamentally does to LLMs. One line of work credits RLVR with unlocking new capabilities in LLMs, pointing to emergent chain‑of‑thought reasoning and strong performance on challenging benchmarks. On the other hand, recent empirical studies (Karan & Du, 2025; Yue et al., 2025; Song et al., 2025) paint a more conservative picture: when given sufficiently large sampling budgets, pre-trained models can often match or even surpass the performance of RL-fine-tuned counterparts. These observations suggest that, rather than endowing models with new skills, RLVR may primarily act to sharpen the output distribution, making the model more confident in its existing knowledge.
@@
We argue that these perspectives are not mutually exclusive. From a distributional perspective, pre‑training and supervised fine‑tuning (SFT) provide the model with a rich, low‑dimensional manifold of knowledge and latent reasoning procedures. RLVR then performs a search over this manifold, guided by verifiable rewards. A moderate degree of sharpening can be beneficial, since it shifts probability mass toward high‑reward semantic modes, making the model more reliable under finite sampling. However, if sharpening is too aggressive, the policy will collapse onto a narrow set of modes, a failure mode we call over‑sharpening. Wherein, the model explores only a small subset of the solution space for each input and many valid modes are discarded. As a result, the training process stops discovering new reasoning strategies, limiting capability growth.
This paper is most closely related to research on entropy collapse. Typical examples include flexible clipping ranges (e.g., DAPO-syple methods (Yu et al., 2025; Wu et al., 2025)), as well as token‑level heuristics that only update high‑entropy tokens (Wang et al., 2025) or truncate gradients on highly confident tokens (Cui et al., 2025). Another family of methods operates at the batch level, discarding degenerate batches (those that are all‑correct or all‑incorrect) and focusing on medium‑difficulty queries, sometimes augmented with carefully constructed synthetic data to keep training in a higher‑entropy regime. Other methods directly add entropy regularization or its variants (Cheng et al., 2025; Prabhudesai et al., 2025) to the RL objective. All of these works share a common idea: entropy is a finite resource that must be prevented from diminishing excessively. However, a comprehensive understanding of how RLVR reshapes the solution space remains unclear. To fill this gap, we study the following critical question:
Can we give a formal account of how RLVR reshapes the solution space and characterize the transition from moderate sharpening to over‑sharpening that destroys diversity and capabilities?
@@
  • Two simple, general‑purpose calibration strategies for RLVR. Building on this analysis, we introduce inverse‑success advantage calibration and distribution‑level calibration. The former one automatically down‑weights updates from easy, high‑success queries and emphasizes harder ones. The later one employ a memory network to reshape the rollout distribution to avoid over‑training on already dominant modes and to promote exploration of alternative reasoning patterns. Both techniques can be plugged into existing RLVR pipelines against over‑sharpening.
  • Extensive experiments over multiple LLMs and benchmark datasets. The performance of the proposed methods over six mathematical reasoning benchmarks consistently outperforms state-of-the-art baselines across both AVG@8 and PASS@8. We also show that our methods are compatible with existing entropy-regularization techniques (e.g., DAPO), delivering additive performance improvements.
@@
The ideal KL‑regularized RLVR objective provably yields moderate sharpening, where the probability of every correct mode is strictly increased (Corollary 2.1).
@@
2 Analyzing Distribution Collapse: Sampling Bias and Semantic Coupling
@@ 2.1 Problem Setup
We consider a policy model $\pi_{\theta}$ initialized from a reference model $\pi_{\text{ref}}$. To facilitate analysis, we abstract the output space of the policy model for a given query $q$ as a set of $K_{1}+K_{2}$ mutually exclusive modes: $K_{1}$ distinct correct modes $\mathcal{O}^{+}=\{o_{1},\dots,o_{K_{1}}\}$ and $K_{2}$ distinct incorrect modes $\mathcal{O}^{-}=\{o_{K_{1}+1},\dots,o_{K_{1}+K_{2}}\}$. We focus on a classic RLVR scenario with binary verifiable rewards where all correct modes share a non‑negative advantage $A_{+}\geq 0$, while all incorrect modes share a negative advantage $A_{-}<0$. Definition 2.1 formalizes two possible behaviours of an updated policy relative to $\pi_{\text{ref}}$.
@@
Intuitively, moderate sharpening reflects a holistic improvement where the model becomes more confident in correct modes without forgetting any valid reasoning patterns. Over‑sharpening instead suppresses some correct modes in favor of others, which could be problematic because the suppressed patterns might be essential for solving other queries. Consequently, such suppression could harm the model’s generalization ability. Besides, to some extent, over-sharpening can serve as a formal characterization of entropy collapse. As suppression intensifies, the policy degenerates into a deterministic distribution. Thus, entropy collapse is the asymptotic limit of severe over-sharpening. In what follows, we first present that infinite‑sample RLVR always yields moderate sharpening, then show how finite‑batch sampling fundamentally alters this behavior and leads to over‑sharpening on other queries via semantic coupling.
@@ 2.2 Sampling Bias
Then, this optimal policy guarantees moderate sharpening:
@@
Suppose $A(q,o_{i})=A_{+}\geq 0$ for $i=1,\cdots,K_{1}$ and $A(q,o_{j})=A_{-}\leq 0$ for $i=K_{1}+1,\cdots,K_{1}+K_{2}$, then Equation 2 yields moderate sharpening.
@@
From Equation 2, for $i\in\mathcal{O}^{+}$, $\pi^{*}(o_{i}|q)=\pi_{\text{ref}}(o_{i}|q)\,e^{A_{+}/\beta}\,/\,Z(q)$. Since $A_{+}\geq 0$, we have $e^{A_{+}/\beta}\geq 1$. Moreover, $Z(q)=\sum_{i\in\mathcal{O}^{+}}\pi_{\text{ref}}(o_{i}|q)e^{A_{+}/\beta}+\sum_{j\in\mathcal{O}^{-}}\pi_{\text{ref}}(o_{j}|q)e^{A_{-}/\beta}\leq e^{A_{+}/\beta}$ because $e^{A_{-}/\beta}\leq 1$. Thus $\pi^{*}(o_{i}|q)\geq\pi_{\text{ref}}(o_{i}|q)$ for $i\in\mathcal{O}^{+}$. ∎
Corollary 2.1 indicates that, in the idealized infinite‑sample regime, Equation 1 never suppresses a correct mode.
Finite‑sample sharpening behavior. In practice, we approximate Equation 1 using Monte Carlo estimation with a finite batch of $G$ independent samples $\{o^{(1)},\dots,o^{(G)}\}\sim\pi_{\theta}(\cdot|q)$. This yields the empirical objective:
$\begin{split}\hat{J}(\theta)=\frac{1}{G}\sum_{s=1}^{G}&A(q,o^{(s)})\log\pi_{\theta}(o^{(s)}|q)\\ &-\beta\,D_{\text{KL}}\bigl(\pi_{\theta}(o^{(s)}|q)\,\|\,\pi_{\text{ref}}(o^{(s)}|q)\bigr).\end{split}$ (3)
While gradient descent updates the policy incrementally, the update direction effectively pushes the model towards the stationary point of Equation 3. That stationary point, which we call the batch-optimal policy $\hat{\pi}$, admits a closed form analogous to Theorem 2.1, but with each advantage weighted by the empirical frequency of the corresponding output in the batch. Let $N_{i}$ denote the number of times mode $o_{i}$ appears in the batch ($\sum_{i}N_{i}=G$). Theorem 2.2 gives the batch‑optimal policy $\hat{\pi}$.
@@
The policy that maximizes Equation 3 is
@@
where $Z^{\prime}(q)=\sum_{j}\pi_{\text{ref}}(o_{j}|q)\exp\!\bigl(N_{j}A(q,o_{j})/(\beta G)\bigr)$.
We emphasize that a single gradient step will generally not reach $\hat{\pi}$ exactly, but the update can be viewed as a geometric interpolation between the current policy and the batch-optimal target (see Theorem B.1), so the qualitative insights carry over.
To understand whether Equation 4 leads to moderate sharpening or over‑sharpening, we examine $\Delta_{i}=\hat{\pi}(o_{i}|q)-\pi_{\text{ref}}(o_{i}|q)$ in which $\sum_{i}\Delta_{i}=0$. We partition the modes ($\mathcal{O}^{+}\cup\mathcal{O}^{-}$) into three categories based on empirical frequency $N_{i}$ and advantage sign $A(q,o_{i})$. We first distinguish between sampled modes ($N_{i}>0$) with positive versus negative advantages. Following the logic of Corollary 2.1, probability mass is naturally reallocated towards sampled modes with positive advantages, while those with negative advantages are suppressed. The key difference lies in the behavior of unsampled modes ($N_{i}=0$).
For any unsampled mode $o_{i}$ with $N_{i}=0$, the exponential term in Equation 4 becomes 1. Consequently, the update simplifies to a uniform scaling:
@@
This equation implies all unsampled modes move in the same direction. If $Z^{\prime}(q)>1$, the probabilities of all unsampled modes decrease, and vice versa. Since the set of unsampled modes could include correct answers, the condition $Z^{\prime}(q)>1$ can lead to over-sharpening at $q$.
@@
Let $S^{+}$ and $S^{-}$ denote the index sets of sampled correct and incorrect modes, respectively. Then the batch partition function $Z^{\prime}(q)$ in Theorem 2.2 satisfies:
@@
Moreover, consider normalized binary advantages $A_{+}=\frac{1-p^{+}}{\sigma},A_{-}=-\frac{p^{+}}{\sigma},$ where $p^{+}$ is the empirical batch accuracy defined by $\sum_{i\in S^{+}}N_{i}=Gp^{+}$, and $\sigma>0$ is the empirical standard deviation. Define $\pi_{\min}^{+}=\min_{i\in S^{+}}\pi_{\mathrm{ref}}(o_{i}\mid q)$, $\pi_{\max}^{-}=\max_{j\in S^{-}}\pi_{\mathrm{ref}}(o_{j}\mid q)$, and $\Delta_{\pi}=\pi_{\min}^{+}-\pi_{\max}^{-}$. The above bound can be simplified to:
$Z^{\prime}(q)\;\geq\;1+\frac{p^{+}(1-p^{+})}{\beta\sigma}\,\Delta_{\pi}.$
See Appendix C for proof.
We now analyze $Z^{\prime}(q)$. As shown in Theorem 2.3, the sharpening behavior is primarily governed by the probability gap $\Delta_{\pi}=\pi_{\min}^{+}-\pi_{\max}^{-}$. Whenever $\Delta_{\pi}>0$, we see $Z^{\prime}(q)>1$ and therefore a uniform suppression of unsampled modes. Intuitively, $\Delta_{\pi}>0$ arises when the model is already performing well on the given query. In fact, this also explains many existing methods (e.g., filtering out high‑confidence/easy samples (Xiong et al., 2025; Amin et al., 2025)) as implicit ways of preventing $Z^{\prime}(q)$ from becoming too large.
Moreover, we observe that the normalization coefficient, $\frac{p^{+}(1-p^{+})}{\sigma}$, acts as a stabilizer by vanishing as $p^{+}\to 0$ or $p^{+}\to 1$, thereby down-weighting updates from highly skewed batches. However, it is insufficient. In other words, the standard normalization indeed can reduce the rate of collapse but fails to address the underlying structural cause, i.e., once the policy becomes confident ($\Delta_{\pi}>0$), the optimization landscape fundamentally favors the suppression of diverse, unvisited modes.
@@ 2.3 Semantic Coupling
We now study how over‑sharpening at a query $q$ affects the model’s behavior at another query $q^{\prime}$. We focus on the case where $q$ and $q^{\prime}$ are positively semantically related, e.g., two math problems that rely on the same underlying theorem. This implies that a parameter update improving the model’s performance on $q$ should also increase the likelihood of the correct response for $q^{\prime}$. We absorb the sign of the advantage $A(q,o)$ into the gradient direction and formalize this in 2.1.
@@
Let $(q,o)$ and $(q^{\prime},o^{\prime})$ be two query–output pairs that are positively semantically related. We assume that their logit gradients are non-negatively aligned:
$\nabla_{\theta}f(q,o)^{\top}\nabla_{\theta}f(q^{\prime},o^{\prime})\;\geq\;0.$
Recall from Theorem 2.2 that the batch‑optimal policy $\hat{\pi}$ for $q$ implies a target logit shift. Let $\mathbf{y}\in\mathbb{R}^{G}$ be the vector of target shifts for the batch $\{o^{(s)}\}_{s=1}^{G}$, where $y_{i}=\frac{N_{i}|A(q,o_{i})|}{\beta G}$ [^1] Due to the sign-absorption convention, the advantage term takes its absolute value here.. The parameter update $\Delta\theta$ induced by fitting this batch (Jacot et al., 2018) is given by [^2] For readers unfamiliar with the optimization dynamics in the linearized regime, we provide the necessary background in Appendix D.:
@@
where $\mathbf{J}_{q}$ is the Jacobian of the logits with respect to parameters, and $\mathbf{K}_{qq}=\mathbf{J}_{q}\mathbf{J}_{q}^{\top}$ is the empirical kernel matrix on the current batch.
For $(q^{\prime},o^{\prime})$, the change in the logit value is induced by the parameter update $\Delta\theta$:
@@
We define $\lambda_{\text{min}},\lambda_{\text{max}}=\min_{i}K_{ii},\max_{i}K_{ii}$ (feature specificity) and $\rho_{\text{min}},\rho_{\text{max}}=\min_{i\neq j}K_{ij},\max_{i\neq j}K_{ij}$ (cross-similarity). We also introduce Assumptions 2.2 and 2.3. First, to render the analysis tractable, we model the shift from a source query to a target query as a signal decay, where gradient correlations are naturally attenuated [^3] In fact, 2.2 can be relaxed to a bounded interval $\eta_{\min},\eta_{\max}]$ for greater generality, which would simply introduce a scaling factor $\frac{\eta_{\min}}{\eta_{\max}}$ to the ratio derived in Corollary 2.2.. Second, we enforce a locality condition, reflecting the intuition that a parameter update intended to reinforce a mode $o$ should boost $o$ itself significantly more than it boosts distinct modes $o^{\prime}$. We then obtain the bound on the induced logit shifts, as shown in Theorem 2.4.
@@
We assume that gradient correlations decay by a factor $\eta\in[0,1]$ when shifting from $q$ to $q^{\prime}$, i.e., $k((q^{\prime},o^{\prime}),(q,o^{(s)}))\approx\eta K((q,o^{\prime}),(q,o^{(s)}))$.
@@
We assume that $\lambda_{\min}\geq\rho_{\max}$, suggesting that the self-alignment consistently exceeds cross-alignment.
@@
See Appendix E for proof.
We now leverage Theorem 2.4 to characterize when over‑sharpening occurs at $q^{\prime}$. Consider a semantically coupled pair $(q,q^{\prime})$ and suppose that at $q^{\prime}$ there exists at least one mode $o^{\prime}$ that appears in the batch for $q$ (Case 2) and at least one mode $\bar{o}$ that does not (Case 1). Over‑sharpening at $q^{\prime}$ arises when the batch‑seen mode receives a much larger logit increase than a semantically related but batch‑unseen mode, so that after normalization the latter is suppressed. Corollary 2.2 makes this comparison explicit.

### v2

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) is a central paradigm for turning large language models (LLMs) into reliable problem solvers, especially in logic‑heavy domains. Despite its empirical success, it remains unclear whether RLVR elicits novel capabilities or merely sharpens the distribution over existing knowledge. We study this by formalizing over-sharpening, a phenomenon where the policy collapses onto limited modes, suppressing valid alternatives. At a high level, we discover finite-batch updates intrinsically bias learning toward sampled modes, triggering a collapse that propagates globally via semantic coupling. To mitigate this, we propose inverse-success advantage calibration to prioritize difficult queries and distribution-level calibration to diversify sampling via a memory network. Extensive evaluations show that our proposed methods can effectively maintain a healthy exploration-exploitation balance and achieve sustained accuracy gains compared to state-of-the-art methods.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a cornerstone approach for specializing large language models (LLMs) to domains where correctness is checkable (DeepSeek-AI, 2025; Yang et al., 2025). However, there is still no clear consensus on what RLVR fundamentally does to LLMs. One line of work credits RLVR with unlocking new capabilities in LLMs, pointing to emergent chain‑of‑thought reasoning and strong performance on challenging benchmarks. On the other hand, recent empirical studies (Karan and Du, 2025; Yue et al., 2025; Song et al., 2025) paint a more conservative picture: when given sufficiently large sampling budgets, pre-trained models can often match or even surpass the performance of RL-fine-tuned counterparts. These observations suggest that, rather than endowing models with new skills, RLVR may primarily act to sharpen the output distribution, making the model more confident in its existing knowledge.
@@
We argue that these perspectives are not mutually exclusive. From a distributional perspective, pre‑training and supervised fine‑tuning (SFT) provide the model with a rich, low‑dimensional manifold of knowledge and latent reasoning procedures. RLVR then performs a search over this manifold, guided by verifiable rewards. We formalize the core mechanism of this process as sharpening and distinguish between two critical regimes: moderate and over-sharpening. Moderate sharpening is beneficial, since it shifts probability mass toward high‑reward semantic modes, making the model more reliable under finite sampling. However, if sharpening is too aggressive, the policy will collapse onto a narrow set of modes, prematurely discarding valid alternative strategies. This restricts the model’s exploration to a small subset of the solution space, eventually stalling the discovery of novel reasoning chains and limiting overall capability growth.
This paper is most closely related to research on entropy collapse, a pervasive phenomenon in RLVR where the policy’s predictive distribution becomes prematurely concentrated. In many respects, over-sharpening serves as a more rigorous and formal characterization of this collapse. Typical examples to mitigate entropy collapse include token-level heuristics like selective updates of high-entropy tokens (Wang et al., 2025) or gradient truncation for highly confident ones (Cui et al., 2025), and batch-level interventions that discard degenerate samples (Xiong et al., 2025) or utilize curated synthetic data (Amin et al., 2025). Additionally, explicit entropy regularization and its variants (e.g., DAPO-style clipping (Yu et al., 2025; Wu et al., 2025) or direct objective penalties (Cheng et al., 2025; Prabhudesai et al., 2025)) are employed. While effective in practice, these methods primarily treat entropy as a finite resource to be preserved through technical mitigations, leaving the underlying mechanisms of over-sharpening largely unexplored. For example, it remains unclear whether over-sharpening is an inevitable byproduct of advancing model performance or a structural artifact of algorithmic bias.
@@
  • Two simple, general‑purpose calibration strategies for RLVR. Building on this analysis, we introduce inverse‑success advantage calibration and distribution‑level calibration. The former one automatically down‑weights updates from easy, high‑success queries and emphasizes harder ones. The later one employs a memory network to reshape the rollout distribution to avoid over‑training on already dominant modes and to promote exploration of alternative reasoning patterns. Both techniques can be plugged into existing RLVR pipelines against over‑sharpening.
[…]
@@
The ideal KL‑regularized RLVR objective is inherently benign (moderate sharpening), where the probability of every correct mode is strictly increased (Corollary 3.1).
@@
2 Related Work
Since the release of OpenAI o1 (Jaech et al., 2024) and DeepSeek-R1 (DeepSeek-AI, 2025), RLVR (or GRPO and its variants) has gained particular traction for enhancing the reasoning ability of LLMs. A key driver of its effectiveness over SFT could be the explicit incorporation of negative samples (Xiong et al., 2025). To stabilize the often volatile training dynamics, a significant body of work focused on refining advantage estimation. Prominent techniques include RLOO (Leave-One-Out) (Kool et al., 2019), DrGRPO (Liu et al., 2025), which modifies the variance term to improve stability, and Reinforce++ with baseline (Hu et al., ) that shifts from sample-level to global batch-level normalization. Moreover, Cui et al. (2025) identified an empirical law showing that performance gains are often traded for entropy, leading to predictable saturation. To prevent such premature convergence, various entropy-management strategies have been proposed. These include token-level heuristics targeting high-entropy forking tokens (Wang et al., 2025), multi-temperature sampling for explicit exploration (Zhuang et al., 2025), and entropy-aware clipping mechanisms (Yu et al., 2025; Wu et al., 2025). RENT (Prabhudesai et al., 2025) and RPN (He et al., 2025) demonstrate that minimizing entropy on confident chains or adding explicit regularization can also drive gains. Despite these advancements, recent studies (Karan and Du, 2025; Yue et al., 2025; Song et al., 2025) indicate that RLVR training can sometimes lead to a degradation in general capabilities. Critically, the underlying mechanics of over-sharpening remain largely unexplored and this paper aims to bridge this gap.
3 Analyzing Distribution Collapse
we first present that infinite‑sample RLVR always yields moderate sharpening, then show how finite‑batch sampling fundamentally alters this behavior and leads to over‑sharpening on other queries via semantic coupling.
@@ 3.1 Problem Setup
We consider a policy model $\pi_{\theta}$ initialized from a reference model $\pi_{\text{ref}}$. For a query $q$, we coarsen the sequence‑level output space into $K_{1}$ mutually exclusive correct modes $\mathcal{O}^{+}=\{o_{1},\dots,o_{K_{1}}\}$ and $K_{2}$ incorrect modes $\mathcal{O}^{-}$. Each mode can be viewed as an equivalence class of outputs that implement the same high‑level reasoning pattern or solution. We focus on the RLVR setting with binary verifiable rewards, where correct modes share a non‑negative advantage $A_{+}\geq 0$, and incorrect modes share $A_{-}<0$. Definition 3.1 formalizes two possible behaviors of an updated policy relative to $\pi_{\text{ref}}$.
@@
Moderate sharpening reflects a holistic improvement where the model gains confidence without forgetting valid reasoning patterns. Over‑sharpening instead suppresses some correct modes in favor of others, which could be problematic because the suppressed patterns might be essential for solving other queries. Moreover, as suppression intensifies, the policy degenerates into a deterministic distribution. Thus, entropy collapse can be viewed as the asymptotic limit of severe over-sharpening.
@@ 3.2 Sampling Bias
Then, as can be seen in Corollary 3.1, Equation 1 guarantees moderate sharpening.
@@
Suppose $A(q,o_{i})=A_{+}\geq 0$ for $o_{i}\in\mathcal{O}^{+}$ and $A(q,o_{j})=A_{-}\leq 0$ for $o_{j}\in\mathcal{O}^{-}$, then Equation 2 yields moderate sharpening.
@@
From Equation 2, for $o_{i}\in\mathcal{O}^{+}$, $\pi^{*}(o_{i}|q)=\pi_{\text{ref}}(o_{i}|q)\,e^{A_{+}/\beta}\,/\,Z(q)$. Since $A_{+}\geq 0$, we have $e^{A_{+}/\beta}\geq 1$. Moreover, $Z(q)=\sum_{o_{i}\in\mathcal{O}^{+}}\pi_{\text{ref}}(o_{i}|q)e^{A_{+}/\beta}+\sum_{o_{j}\in\mathcal{O}^{-}}\pi_{\text{ref}}(o_{j}|q)e^{A_{-}/\beta}\leq e^{A_{+}/\beta}$ because $e^{A_{-}/\beta}\leq 1$. Thus $\pi^{*}(o_{i}|q)\geq\pi_{\text{ref}}(o_{i}|q)$ for $o_{i}\in\mathcal{O}^{+}$. ∎
Finite‑sample sharpening behavior. In practice, we approximate Equation 1 using a finite batch of $G$ samples $\{o^{(1)},\cdots,o^{(G)}\}\sim\pi_{\theta}(\cdot|q)$. This yields an empirical objective where the gradient updates incrementally push the policy toward its stationary point. We characterize this stationary point as the batch-optimal policy $\hat{\pi}$, which admits a closed-form solution analogous to the exact case but incorporates the empirical frequency $N_{i}$ of each mode in the batch. Let $N_{i}$ denote the number of times mode $o_{i}$ appears in the batch ($\sum_{i}N_{i}=G$). Theorem 3.2 gives the batch‑optimal policy $\hat{\pi}$.
@@
The batch-optimal policy $\hat{\pi}$ is given by
@@
where $Z^{\prime}(q)=\sum_{j}\pi_{\text{ref}}(o_{j}|q)\exp\!\bigl(N_{j}A(q,o_{j})/(\beta G)\bigr)$. See Appendix B for proof.
We emphasize that a single gradient step will generally not reach $\hat{\pi}$ exactly, but the update can be viewed as a geometric interpolation between the current policy and the batch-optimal target (see Theorem C.1), so the qualitative insights carry over.
To understand the sharpening dynamics of Equation 3, we examine $\Delta_{i}=\hat{\pi}(o_{i}|q)-\pi_{\text{ref}}(o_{i}|q)$ in which $\sum_{i}\Delta_{i}=0$. We partition the modes ($\mathcal{O}^{+}\cup\mathcal{O}^{-}$) into three categories based on empirical frequency $N_{i}$ and advantage sign $A(q,o_{i})$. Following the logic of Corollary 3.1, sampled modes ($N_{i}>0$) reallocate mass according to advantage signs. For any unsampled mode $o_{i}$ with $N_{i}=0$, the exponential term in Equation 3 becomes 1. Consequently, the update simplifies to a uniform scaling:
@@
If $Z^{\prime}(q)>1$, the probabilities of all unsampled modes decrease, and vice versa. Since this set may contain valid reasoning paths, $Z^{\prime}(q)>1$ can induce over-sharpening.
@@
(Proof in Appendix D.) Let $S^{+}$ and $S^{-}$ be the indices of sampled correct and incorrect modes. Then the batch partition function $Z^{\prime}(q)$ in Theorem 3.2 satisfies:
@@
For normalized binary advantages $A_{+}=(1-p^{+})/\sigma,A_{-}=-p^{+}/\sigma$ with batch accuracy $p^{+}$ and standard deviation $\sigma$, this bound simplifies to: $Z^{\prime}(q)\;\geq\;1+\frac{p^{+}(1-p^{+})}{\beta\sigma}\,\Delta_{\pi}.$ where $\Delta_{\pi}=\min_{i\in S^{+}}\pi_{\mathrm{ref}}(o_{i}|q)-\max_{j\in S^{-}}\pi_{\mathrm{ref}}(o_{j}|q)$.
As shown in Theorem 3.3, the sharpening behavior is governed by the probability gap $\Delta_{\pi}$. Whenever $\Delta_{\pi}>0$, we see $Z^{\prime}(q)>1$ and therefore a uniform suppression of unsampled modes. Intuitively, $\Delta_{\pi}>0$ arises when the model is already performing well on the given query. In fact, this also explains many existing methods (e.g., filtering out high‑confidence/easy samples (Yu et al., 2025; Xiong et al., 2025; Amin et al., 2025; Bae et al., 2025)) as implicit ways of preventing $Z^{\prime}(q)$ from becoming too large. Furthermore, while the term $p^{+}(1-p^{+})/\sigma$ stabilizes updates from skewed batches (vanishing as $p^{+}\to\{0,1\}$), it is insufficient to counteract the intrinsic optimization bias that favors suppressing unsampled modes once model confidence is established.
@@ 3.3 Semantic Coupling
We now analyze how over‑sharpening propagates from a query $q$ to a positively semantically related query $q^{\prime}$, e.g., two math problems that rely on the same underlying theorem. This implies that a parameter update improving the model’s performance on $q$ should also increase the likelihood of the correct response for $q^{\prime}$. To formalize this, we absorb the sign of the advantage $A(q,o)$ into the gradient direction and adopt 3.1.
@@
For positively semantically related pairs $(q,o)$ and $(q^{\prime},o^{\prime})$, their logit gradients are non-negatively aligned: $\nabla_{\theta}f(q,o)^{\top}\nabla_{\theta}f(q^{\prime},o^{\prime})\;\geq\;0.$
The batch-optimal policy $\hat{\pi}$ prescribes target logit shifts $\mathbf{y}\in\mathbb{R}^{G}$, where $y_{i}=\frac{N_{i}|A(q,o_{i})|}{\beta G}$ [^1] Due to the sign-absorption convention, the advantage term takes its absolute value here.. With a sufficiently small learning rate (Jacot et al., 2018), fitting these targets yields the parameter update [^2] For readers unfamiliar with the optimization dynamics in the linearized regime, we provide the necessary background in Appendix E.:
@@
where $\mathbf{J}_{q}$ is the Jacobian of the logits with respect to parameters, and $\mathbf{K}_{qq}=\mathbf{J}_{q}\mathbf{J}_{q}^{\top}$ is the empirical kernel matrix on the current batch. The update induces a logit shift at $(q^{\prime},o^{\prime})$:
@@
Let $\mathbf{M}(\lambda,\rho)=(\lambda-\rho)\mathbf{I}+\rho\mathbf{1}\mathbf{1}^{\top}$, which possesses a principal eigenvalue $\lambda+(G-1)\rho$ associated with the constant vector $\mathbf{1}$, and a (G1)(G-1)-degenerate eigenvalue $\lambda-\rho$ for the remaining subspace. Then, we can always find tight spectral bounds to encapsulate the actual spectrum of $\mathbf{K}_{qq}$: $0\preceq\mathbf{M}(\lambda_{\min},\rho_{\min})\preceq\mathbf{K}_{qq}\preceq\mathbf{M}(\lambda_{\max},\rho_{\max}).$ Here, $\lambda_{\min}$ and $\lambda_{\max}$ bound the diagonal entries (self‑alignment of each mode), while $\rho_{\min}$ and $\rho_{\max}$ bound the off‑diagonal entries (cross‑alignment between different modes). In typical scenarios, updates meant for a particular mode influence that mode more than any other, which translates to $\rho_{\min}\leq\rho_{\max}<\lambda_{\min}\leq\lambda_{\max}$. We formalize this as 3.3. See Appendix F for empirical validation of 3.1 and 3.3. Then, to make the analysis manageable, we make 3.2 that models the shift from a source query to a target query as a signal decay3333.2 can be relaxed to a bounded interval $\eta_{\min},\eta_{\max}]$ for greater generality, which would simply introduce a scaling factor $\frac{\eta_{\min}}{\eta_{\max}}$ to the ratio derived in Corollary 3.2.. Intuitively, $\eta$ can be interpreted as the similarity between $q^{\prime}$ and $q$ from the model’s perspective. We then obtain the bound on the induced logit shifts, as shown in Theorem 3.4.
@@
Gradient correlations decay by a factor $\eta\in[0,1]$ when shifting from query $q$ to $q^{\prime}$, modeled as $k((q^{\prime},o^{\prime}),(q,o^{(s)}))\approx\eta K((q,o^{\prime}),(q,o^{(s)}))$.
@@
Self-alignment consistently exceeds cross-alignment: $\rho_{\min}\leq\rho_{\max}<\lambda_{\min}\leq\lambda_{\max}.$
@@
See Appendix G for proof.
We leverage Theorem 3.4 to characterize when over‑sharpening occurs at $q^{\prime}$. Consider a semantically coupled pair $(q,q^{\prime})$ and suppose that at $q^{\prime}$ there exists at least one mode $o^{\prime}$ that appears in the batch (Case 2) and at least one mode $\bar{o}$ that does not (Case 1). Over‑sharpening at $q^{\prime}$ occurs when the logit increase for the batch‑seen mode dominates that for the batch‑unseen mode, because after normalization the latter becomes suppressed. To quantify this, Corollary 3.2 compare the guaranteed minimum increase for a seen mode (the lower bound of Case 2) to the maximum possible increase for an unseen mode (the upper bound of Case 1).